Radial nodes present in 3s and 3p-orbitals are respectively

Grade:12

1 Answers

Ravneet
5 years ago
radial nodes: n – l – 1 
where n principal quantum numbers (represent size and energy )
           l azimuthal quantum number  ( represent shape )
for s, the value of l = 0 
for p, the value of p = 1 
 
now 3s, radial node = 3-0-1 = 2 
        3p, radial node = 3-1-1 = 1